Write a balanced equation for 106/46 Of (a,p) 109/47 Ag
1
Expert's answer
2021-07-17T04:52:10-0400

"_{46}^{106}Pd(\alpha, P) _{47}^{109} Ag \\\\
 _2^4 \alpha=_2^4 He \space Nuclei \space \& \space P=proton\\\\
_{46}^{106}Pd +_2^4 \alpha \to _{47}^{109} Ag +P"